OpenStudy (anonymous):

Find the zeros of each function by factoring f(x)=x^2-15x+54

OpenStudy (anonymous):

f(x)=(x+6)(x+9) zeros are x=-6, x=-9

OpenStudy (anonymous):

ooops f(x)=(x-6)(x-9) x=6, x=9